OpenStudy (anonymous):

∑[n=1,∞,(15-x^(2))^(n)]

OpenStudy (anonymous):

well here is the answer, lets see if we can get there http://www.wolframalpha.com/input/?i=sum+n+%3D+1+to+infinity+%2815-x^2%29^n

OpenStudy (jamesj):

Remember first that for |r| < 1, \[ 1 + r + r^2 + ... + r^n + r^{n+1} + ... = \frac{1}{1-r} \] Now use that relation with your expression above. I.e., r = 15 - x^2

OpenStudy (anonymous):

think it might be a simple as using \[\sum x^n=\frac{1}{1-x}\] with some adjustments

OpenStudy (anonymous):

what jamesj said

OpenStudy (jamesj):

Deriving the formula for the sum is pretty straight-forward. What requires a few more neurons is determining for what values of x the infinite sum exists.

OpenStudy (anonymous):

well i have a question. why the "-1" i probably would have missed that one

OpenStudy (jamesj):

because in the 'geometric series' of the question, there is no "1" term in the sum. It starts with r; not 1.

OpenStudy (anonymous):

right. i often mess that one up, unless i start from scratch. as for radius of convergence, it is pretty damned small i think

OpenStudy (jamesj):

It's the x such that \[ |15-x^2| < 1 \] It's a more interesting set than most convergence intervals.

OpenStudy (jamesj):

@charitha, you following?

OpenStudy (anonymous):

@JamesJ yes,, thanks :)

OpenStudy (anonymous):

oh there are two intervals! is this part of the question?

OpenStudy (anonymous):

@satellite73 ya